Your firm currently has 1,730,483 shares of stock outstanding that sell for 16.02. Additionally, you've issued 24,166 bonds, with a par value of S1,000. They currently sell for 85.6% of par.

What is the firm's debt weight?

(Enter your answer as a percentage to 1 decimal place. ex: 12.3)
Answer #1

Total value of:

Equity=(1,730,483*16.02)=$27722337.7

Bonds=24,166*(1000*85.6%)=$20686096

Total value=27722337.7+20686096=$48408433.7

Hence weight of debt=20686096/48408433.7

=42.7%(Approx).